Dear teacher,
I have here an equation algebra question:- 5(-5p+2)= -3(3p+5)

Found 2 solutions by Alan3354, Theo:
Answer by Alan3354(69443)

- 5(-5p+2)= -3(3p+5)
--------
5(-5p+2)= 3(3p+5)
-25p + 10 = 9p + 15
10 = 34p + 15
-5 = 34p
p = -5/34

Answer by Theo(13342)

equation is:

-5(-5p+2)= -3(3p+5)
remove parentheses by performing the multiplication (distributive law of multiplication applies) to get:
25p - 10 = -9p - 15
add 9p to both sides of the equation to get:
34p - 10 = -15
add 10 to both sides of the equation to get:
34p = -5
divide both sides of the equation by 34 to get:
p = -5/34

the answer is confirmed to be good by replacing p with -5/34 and going through the calculations (i used a calculator).